2 pendulums of lenghts 1.44m and 1m start to swing together. The number of vibrations after which they will again start swinging together is.....

Grade:12

1 Answers

Vikas TU
6 years ago
The day and age is straightforwardly relative to the square base of the length of pendulum 
T α √L 
T1/T2 = √L1/L2 
= √1.44/1 
= 12/10 
10T1 = 12T2 
After 10 motions they will swing together.